OpenStudy (anonymous):

1-2sin^2x=0 Solve the equation in the interval 0

hartnn (hartnn):

can you find 2 values of sin x first ?

OpenStudy (anonymous):

sin x = sqrt(1/2) If you look at the trigonometric table, you will see that sin 45 = sqrt(1/2) Hence 0<x<360 gives you the answers x = 45, (90+45 = 135) pi/4 radians, 3/4 pi radians.